14. The normal to the curve y = X^2– 4x at the point ( 3 , - 3 ) cuts the X-axis at A and the

Y-axis at B. Find the equation of the normal and the coordinates of A and B.


1
Expert's answer
2021-09-20T06:26:21-0400
"y=x^2-4x"

"y'=(x^2-4x)'=2x-4"

The slope of the tangent to the curveat the point "(3, -3)" is


"slope_1=m_1=y'(3)=2(3)-4=2"

The slope of the normal to the curve at the point "(3, -3)" is


"slope_2-m_2=-\dfrac{1}{m_1}=-\dfrac{1}{2}"

The equation of the normal to the curve at the point "(3, -3)" is


"y-(-3)=-\dfrac{1}{2}(x-3)"

The equation of the normal to the curve at the point "(3, -3)" in slope-intercept form is

"y=-\dfrac{1}{2}x-\dfrac{3}{2}"

"x" -intercept: "y=0=>-\dfrac{1}{2}x-\dfrac{3}{2}=0=>x=-3"

"A(-3, 0)"


"y" -intercept: "x=0=>y=-\dfrac{1}{2}(0)-\dfrac{3}{2}=>y=-\dfrac{3}{2}"

"B(0, -\dfrac{3}{2})"